OpenStudy (anonymous):

Can some one help me with this math problem. Select ALL of the equivalent equations. a. 1.2 = 1.8 + m b. 1.2 – m = 1.8 c. 1.8 – 1.2 = m d. m + 1.2 = 1.8 e. 1.2 – 1.8 = m f. m = 1.2 + 1.8 please respond before 11/6/15

OpenStudy (koikkara):

@cinnabun23 ... You have asked for help, not answer ! So please answer the way you know, and we will help you out..

OpenStudy (anonymous):

i don't want an answer just help i don't quite get it

Directrix (directrix):

I don't know that this is a complete definition for equivalent equations but it will suffice for this problem, I think. Two equations are said to be equivalent if they have exactly the same solutions. Sample Problems •The equations x = 1 and 2x = 2 are equivalent, since the only solution to either equation is 1. •The equations 2x = 0 and 3x = 0 are equivalent, since the only solution to either equation is 0. •The equations x = 1 and x = 2 are not equivalent, since their solutions, 1 and 2, respectively, differ.

Directrix (directrix):

I suggest that you solve each of the given equations for m and look for those equations which have the same solution. To help you get started: a. 1.2 = 1.8 + m Subtract 1.8 from both sides 1.2 - 1.8 = m - .6 = m b. 1.2 – m = 1.8 Subtract 1.2 from both sides -m = 1.8 - 1.2 -m = .6 Multiply both sides by 1 m = -.6 Therefore, equations A and B are equivalent. That leaves four more for you to solve.

OpenStudy (anonymous):

thank you so much for helping me :)